OpenStudy (anonymous):

Hello, In the 2nd experiment of lec.4, why would the bullet hit the monkey independent from the magnitude of the initial velociy? thanks

OpenStudy (mos):

time needed to reach the monkey: t=d/u grop of monkey: h1=1/2 g t drop of bulet :h2=1/2 g t h1=h2

OpenStudy (anonymous):

to make it easier as a first approximation: when the velocity is faster it will reach the line where the monkey falls down sooner and higher. the monkey falls with the same velocity every time - therefore it will not fall as deep as it would fall if the velocity of the bullet's lower. keep in mind that the arch is getting wider when the bullet's velocity is higher - therefore it will reach the rod where the monkey hangs on a higher level than if the bullet is slow (as the video sows the bullet can reach the rod at its lowest point or not at all, too).

OpenStudy (anonymous):

Think about it this way: 1. You are aiming the gun directly at the monkey. If we forget about gravity (so we say that the monkey doesn't drop from the tree and the golf ball goes in a straight line as opposed to a parabola) the ball will obviously just hit the monkey. No matter how much faster we shoot the golf ball, it will still hit the monkey, albeit a bit sooner. 2. Now bring gravity back in. As soon as you pull the trigger, the bullet begins it's flight through the air and gravity pulls the golf ball down towards the ground. Also when you pull the trigger, the monkey let's go from the top of the tree and gravity begins to pull the monkey to the ground. Remember that gravity doesn't care about mass of the object, so whether it's a monkey, an apple, a bowling ball, or a golf ball, they will all fall at 9.81 m/s^2. In other words, gravity is pulling both the golf ball and the monkey out of the tree at the same time and to the same degree. The golf ball and monkey fall in unison. The key is to realize that since they are falling in unison, even though to our eyes it looks like the path of the golf ball is a parabola, to the golf ball itself (if it could see and think) it thinks it's always pointed directly at the monkey, because even though the golf ball falls, the monkey falls with it, and so the monkey is always "in the golf balls sites".

OpenStudy (anonymous):

thanks for your help guys.